### v1

# R2GenCSR: Retrieving Context Samples for Large Language Model based X-ray Medical Report Generation

Abstract
Inspired by the tremendous success of Large Language Models (LLMs), existing X-ray medical report generation methods attempt to leverage large models to achieve better performance. They usually adopt a Transformer to extract the visual features of a given X-ray image, and then, feed them into the LLM for text generation. How to extract more effective information for the LLMs to help them improve final results is an urgent problem that needs to be solved. Additionally, the use of visual Transformer models also brings high computational complexity. To address these issues, this paper proposes a novel context-guided efficient X-ray medical report generation framework. Specifically, we introduce the Mamba as the vision backbone with linear complexity, and the performance obtained is comparable to that of the strong Transformer model. More importantly, we perform context retrieval from the training set for samples within each mini-batch during the training phase, utilizing both positively and negatively related samples to enhance feature representation and discriminative learning. Subsequently, we feed the vision tokens, context information, and prompt statements to invoke the LLM for generating high-quality medical reports. Extensive experiments on three X-ray report generation datasets (i.e., IU-Xray, MIMIC-CXR, CheXpert Plus) fully validated the effectiveness of our proposed model. The source code of this work will be released on https://github.com/Event-AHU/Medical_Image_Analysis.
1 Introduction
X-ray image based medical report generation is one of the typical applications of Artificial Intelligence (AI) in healthcare. It aims to utilize powerful AI models to directly generate high-quality medical reports from given X-ray images, thereby alleviating the workload of doctors and reducing patient waiting times. Although the performance of this task has made significant strides with the development of AI, it still falls short of matching the expertise of professional physicians due to various challenges. Due to privacy concerns surrounding X-ray data, there is a lack of quality and diversity in the training datasets, and the rarity of certain diseases and abnormal conditions also results in poor generalization performance of the models. Therefore, the research on X-ray image based report generation is still a very important but unsolved problem.
Figure 1: Comparison between (a-c). existing X-ray report generation frameworks and (d). our newly proposed one; (e). t-SNE feature distribution of our sampled positive and negative context samples from IU-Xray dataset.
Due to the good generalization of deep learning, the performance of X-ray medical report generation has seen steady improvements. For example, Cao et al. propose the Multi-modal Memory Transformer Network (MMTN) [5] for image-report consistent medical report generation. Jin et al. introduce the PromptMRG [19] which attempts to enhance the diagnostic accuracy in the report by feeding diagnosis-aware prompts. Li et al. proposed a new radiological reports DCL [24] framework that uses dynamic graphs to integrate specific and general knowledge to improve visual representation. This framework also enhances visual and textual representation by leveraging contrastive learning objectives. KiUT [18] utilizes the U connection between the encoder and decoder to effectively integrate different levels of visual information. It introduces a knowledge graph and uses a distillation technique to produce reports that are more aligned with real-world conditions. METransformer [46] simulates multi-expert joint diagnostics by introducing multiple learnable “expert” tokens. These tokens focus on different areas while interacting with each other to capture reliable and complementary visual information, facilitating the parallel generation of diagnostic reports. Although these models have achieved good results, their overall performance is still far from reaching the level of human experts. How to narrow the gap further with professional doctors remains a question worth pondering and researching.
Inspired by the great success of LLMs in natural language processing (NLP), some researchers also introduced LLMs for medical report generation. Specifically, Liu et al. [26] propose to augment the LLMs-based radiology report generation using in-domain instance induction and coarse-to-fine decoding. Although better performance can be obtained, these models still achieve their efficiency and performance bottlenecks due to the following issues: Firstly, the performance of current LLMs heavily relies on the tokens humans feed into, for example, the prompt sentence, visual tokens, etc. More comprehensive inputs can guide the LLMs to generate high-quality medical reports. However, seldom of current works consider the context samples (e.g., the samples with/without disease) which may be very important cues for the text generation of the current sample. Secondly, they adopt the Transformer network as the vision backbone which is computationally expensive ($\mathcal{O}(N^{2})$). When handling long-range visual tokens (e.g., the raw X-ray images are usually high-definition), the self-attention in the Transformer often performs poorly on the speed, memory usage, etc.
With these issues in mind, in this work, we propose a novel X-ray medical report generation framework that adopts the Mamba [14] as the backbone and mines the context samples to guide the large language models for high-quality report generation. A comparison between existing models and ours is illustrated in Fig. 1 (a-d). We partition the input X-ray image into patches and project them into visual tokens. Then, the Mamba backbone network is adopted for efficient and effective visual token processing. More importantly, we retrieve some context samples from the training subset for each image in the mini-batch to help our report generator understand which samples are likely to have diseases and which do not. These context samples are also processed using the Mamba backbone and subtracting the global tokens of the current image to get the residual tokens. We provide the context prompts to assist the LLM in distinguishing whether certain tokens are related to diseases. As shown in Fig. 1 (e), it is easy to find that these retrieved positive and negative samples are easily distinguishable from the perspective of the t-SNE [40] feature distribution. Finally, we feed these visual tokens, context tokens, and prompts into the LLM for high-quality X-ray medical report generation. Extensive experiments on two widely used benchmark datasets fully validated the effectiveness of our proposed framework.
To sum up, the main contributions of this work can be listed as follows:
1). We propose a novel large language model based X-ray medical report generation framework that is augmented by context samples in the training phase, termed R2GenCSR.
2). We propose an efficient and effective half-precision vision Mamba that achieves comparable performance to the widely used Transformer backbone network for the X-ray medical report generation task.
3). Extensive experiments on the widely used IU-Xray, MIMIC-CXR, and CheXpert Plus datasets fully validated the effectiveness of our proposed X-ray report generation framework.
2 Related Work
In this section, we will review the related works on the X-ray Medical Report Generation, Large Language Models, State Space Models, and Context Sample Retrieval. More related works can be found in the following surveys [44, 53].
2.1 X-ray Medical Report Generation
Existing X-ray medical report generation models can be divided into CNN (Convolutional Neural Networks)-based, RNN (Recurrent Neural Networks)-based, and Transformer-based frameworks. To be specific, Li et al. [23] propose a model that combines CNNs and RNNs to generate medical reports from chest X-ray images. Jing et al. [20] develop a medical report generation model based on the LSTM [16] (Long Short-Term Memory) framework. They first predict the possible diseases using LSTM and then generate medical reports based on those predictions. Chen et al. [7] demonstrate the effectiveness of generating detailed and accurate radiological reports from X-ray images using the Transformer model, leveraging visual and text data to improve performance. Wang et al. [43] pre-train a ViT model on the high-resolution X-ray images using masked auto-encoder for medical report generation.
2.2 Large Language Models
The combination of medical report generation and Large Language Models (LLMs) has become the most popular research direction in report generation. LLMs are the focus of current research which can be divided into two categories: base LLMs and professional LLMs. The most famous early LLM is Google’s BERT [11], which understands text via a bidirectional encoder, significantly improving the performance of natural language processing tasks by pre-training on large amounts of unlabeled text and then fine-tuning on specific tasks. Meta develops a foundational large language model, LLaMA [37], with several billion to several hundred billion parameters. LLaMA significantly reduces computing resources and energy requirements while maintaining high performance, demonstrating broad potential in practical applications. GPT [30] (Generative Pre-trained Transformer) series consists of large-scale language models developed by OpenAI for natural language generation and understanding. GPT-3 [4] is renowned for its 175 billion parameters and its powerful capabilities in both generating and understanding language.
For the LLMs developed for medical domains, R2Gen-GPT [45] proposes a medical report generation method based on LLM which combines the image and text and fed into the decoder Llama2-7B [38] for report generation. RGRG [36] applies a practice similar to object detection tasks to medical report generation by using GPT-2 [32] to generate separate sentences for the detected areas and then reconnect the sentences. MedicalGPT [48] is a healthcare language modeling project based on the ChatGPT training process. Developed by Google, Med-PaLM2 [33] is a medical language model that combines an improved foundational language model (PaLM 2 [1]), domain-specific fine-tuning for the medical field. Med-PaLM2 improves performance by over 19% compared to Med-PaLM [39] and has reached a new state-of-the-art level. MedVersa [54] is capable of handling multimodal medical inputs and outputs, supporting real-time task specification. Inspired by these work, we propose to further improve the quality of X-ray medical reports via large language models in this paper.
2.3 State Space Model
Due to the high computational cost in the widely used Transformer networks, the State Space Model (SSM) [44] is proposed to achieve linear complexity. To be specific, the S4 [15] model introduces a new structured state space approach through a layered and modular design. This improves the modeling ability of long sequence dependencies and significantly enhances the efficiency and accuracy of sequence modeling. S5 [35] is an in-depth improvement and simplification of the S4 model, aiming to enhance computational efficiency and ease of use while retaining powerful sequence modeling capabilities. Mamba [14] proposes a time-varying state-space model based on a selection mechanism to efficiently model long sequences. With the success of Mamba, researchers have applied it to a variety of research fields. In the field of vision related to medical report generation, the SSM-only approach Vim [55] (Vision Mamba) has achieved good results in terms of performance and efficiency, especially for processing high-resolution images. This is accomplished by adding positional embeddings to image sequences and utilizing bi-directional SSMs to compress visual representations. VMamba [28] proposes a visual Mamba model with a global receptive field and linear complexity. Its success comes from the Cross-Scan Module (CSM), which scans simultaneously from all four corners of the feature map, ensuring that each element in the feature map integrates information from all other locations in different directions. Mamba-2 [9] is an improved version based on the Mamba architecture. By incorporating SSD theory and structural attention mechanisms, enhances performance and efficiency while maintaining the advantages of Mamba. Inspired by the linear computational cost, in this work, we propose to encode the X-ray image using a half-precision vision Mamba network and achieve similar performance on three X-ray medical report generation benchmark datasets.
2.4 Context Sample Retrieving
Retrieval-Augmented Generation (RAG) is a hybrid approach that combines retrieval and generation to enhance the performance of NLP tasks, particularly those requiring extensive knowledge and contextual information. BGFormer [42] (Batch-Graph Transformer) introduced a new Transformer architecture called SSA (Structure-constrained Self-attention), which deeply mines the relationships between samples to provide a new method for robust and differentiated data representation and learning. CricaVPR [29] generates more robust image features by correlating multiple images in a batch through an attention mechanism, using cross-image differences like perspective and illumination as cues for feature learning. CRAG [50] introduces a search evaluator to assess the quality of retrieved documents and enhance search results through large-scale web searches. At the core of EgoInstructor [47] is the retrieval-augmented module, which utilizes existing third-person video resources to help the model better understand and describe first-person perspective video content. Retrieval augmentation can significantly improve generation quality and reduce dependence on the size of the training dataset. RALF [17] (Retrieval-Augmented Layout Transformer), proposed by Horita et al., enhances the generation process by retrieving layout examples most similar to the input image, thereby overcoming the challenge existing methods face in capturing high-dimensional layout structures when training data is limited. EVCAP [22] is a retrieval-augmented image captioning method based on external visual-name memory. It constructs external memory using object images and names, and generates image captions through a retrieval-augmented model. In the field of medical report generation, RAG can also serve as a clinical decision support tool by combining medical databases and research papers, helping physicians quickly access the latest research on disease diagnosis, treatment options, and drug information. Our extensive experiments on three benchmark datasets support the effectiveness of context samples for the medical report generation task.
Figure 2: An overview of our proposed context sample augmented large language model for efficient X-ray medical report generation, termed R2GenCSR. Three main modules are involved in this framework, including the Mamba vision backbone, context sample retrieval, and large language model (LLM). We first extract the visual tokens of the input X-ray image using the Mamba backbone, then, retrieve context samples from the training subset. We get the residual tokens by subtracting the tokens of the input image and its context samples. The LLM takes the vision tokens, context residual tokens, and prompt statements as input and generates a high-quality medical report.
3 Methodology
In this section, we will first give a review of the Mamba network and an overview of our proposed R2GenCSR framework. Then, we will dive into the details of the R2GenCSR framework, with a focus on Input Representation, Context Sample Retrieval, LLM for Report Generation, and Loss Function. More details will be introduced in the subsequent subsections.
3.1 Preliminary: Mamba
Current widely used Mamba networks are developed based on the continuous State Space Model (SSM). It maps a one-dimensional function or sequence $x(t)\in\mathbb{R}^{p}$ to $y(t)\in\mathbb{R}^{q}$ through a hidden state $h(t)\in\mathbb{R}^{n}$. The computing procedure can be summarized as follows:
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (1)
$y(t)$ $=\mathbf{C}h(t),$ (2)
where $\mathbf{A}\in\mathbb{R}^{n\times n}$, $\mathbf{B}\in\mathbb{R}^{n\times p}$, $\mathbf{C}\in\mathbb{R}^{q\times n}$ denotes the state matrix, input matrix, and output matrix.
As the image and text we processed are discrete data, the aforementioned continuous SSMs needed to be transformed into discrete ones. For example, the S4 [15] and Mamba model adopts the Zero-Order Hold (ZOH) to realize this goal, i.e.,
$\bar{\mathbf{A}}$ $=\exp(\Delta\mathbf{A}),$ (3)
$\bar{\mathbf{B}}$ $=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}.$ (4)
where the $\Delta$ is a timescale parameter (also called step size). Thus, we can reformulate the discrete version of SSM as:
$h_{t}$ $=\bar{\mathbf{A}}h_{t-1}+\bar{\mathbf{B}}x_{t},$ (5)
$y_{t}$ $=\mathbf{C}h_{t}.$ (6)
To further strengthen the SSM, Gu et al. propose the Mamba [14] which makes the model varying from time-invariant to dependent. And also speed the training and inference using a couple of hardware-aware algorithms. Inspired by the success of Mamba in natural language processing, researchers also adapt it to the computer vision community, e.g., the VMamaba [28] used in this paper, and vision Mamba [55]. We prefer the readers to check the reference [44] for more details.
3.2 Overview
In this paper, we propose a novel contextual sample retrieval guided large language model framework for efficient X-ray medical report generation. As shown in Fig. 2, it can be divided into three main parts, i.e., the Mamba vision backbone, context retrieval module, and large language model (LLM) for report generation. Given the X-ray image, we first extract its visual tokens using the Mamba backbone. Meanwhile, we retrieve context samples (X-ray samples with and without disease) from the training subset based on the input image and embed them into visual and text tokens. Then, the residual tokens which measure the difference between the input and context samples can be obtained via the subtract operator. Finally, we feed the vision tokens, context residual tokens, and prompt statements into the LLM to generate a high-quality medical report. One can note that the proposed framework stands out from existing methods by incorporating context retrieval and using a linear complexity vision backbone, which enhances feature representation and discriminative learning while maintaining computational efficiency.
3.3 R2GenCSR Framework
In this subsection, we will introduce the R2GenCSR framework from the perspective of Input Representation, Context Sample Retrieval, LLM for Report Generation, and Loss Function.
3.3.1 Input Representation
Assume the dataset contains $\mathbf{N}$ X-ray images, $\mathbf{X}=\{x_{1},x_{2},\dotsc,x_{N}\}$, where each $x_{i}\in\mathbb{R}^{C\times H\times W}$ represents a X-ray image with $C$ channels, height $H$, and width $W$. For each X-ray image $x$, the corresponding feature map $v\in\mathbb{R}^{\hat{H}\times\hat{W}\times\hat{C}}=VMamba(x)$ can be obtained after feeding it into the VMamba backbone, where $\hat{H}$ and $\hat{W}$ are the spatial dimensions of the feature map, and $\hat{C}$ is the number of feature channels. The reason our framework adopts VMamba instead of conventional visual Transformer models, such as ViT and Swin-Transformer, is because the computational complexity of this model is linear ($\mathcal{O}(N)$), requiring lower computational resources. As shown in Fig. 2, the basic VMamba block consists of Layer Normalization (LN), Linear Layer, DW-Conv layer, SiLU activation layer, SS2D module, and also the skip connections.
Then, two distinct types of representations are generated based on feature map $v$, i.e., global features $v_{G}$ and sequential tokens $v_{S}$. Specifically, a 2D global average pooling is applied to $v$ over the spatial dimensions to get the global features $v_{G}\in\mathbb{R}^{\hat{C}}$. The sequential tokens $v_{S}\in\mathbb{R}^{\hat{H}\hat{W}\times E}$ are obtained by flattening the $v$ along the spatial dimension, then, processed by projection layer (Proj) and layer norm (LN) operations. Mathematically speaking,
$v_{G}$ $=avgPool2D(v)$ (7)
$v_{S}$ $=LN(Proj(Flatten(v)))$ (8)
Here, global features $v_{G}\in\mathbb{R}^{\hat{C}}$ capture the global information of the X-ray image, meanwhile, the sequential tokens $v_{S}\in\mathbb{R}^{\hat{H}\hat{W}\times E}$ learns the channel representations.
3.3.2 Context Sample Retrieval
Given the visual tokens of the input X-ray images, one can directly feed them into the large language model to generate the medical reports, clearly, this approach will still reach its performance bottleneck. The context learning methods [42, 12] suggest that other samples in the training set may still play a positive role in each round of model optimization. Therefore, we consider designing a method to mine samples that are positively and negatively correlated with the current sample to enhance discriminative feature learning, thereby better guiding the LLM to generate accurate medical reports.
$\bullet$ Positive and Negative Sample Selection.  For each sample in a mini-batch, we can retrieve its context samples based on keywords in the medical reports. In our implementation, we exploit two different approaches: 1). We adopt the CheXbert [34] to find the possible 14 kinds of diseases from the annotated medical report. If the sample is annotated as No Finding, we treat it as a negative context sample (without disease), otherwise positive (with disease). 2). We find that the medical report with and without Note symbol can be roughly divided based on visual features, as illustrated in Fig. 1 (e). Similarly, we can treat the context samples with/without Note as the positive/negative samples, respectively. In addition, we also randomly select context samples to augment the training of our proposed R2GenCSR model.
After the context samples are retrieved, we extract the global visual features of X-ray images as $\{v_{g,1},v_{g,2},\dotsc,v_{g,n}\},v_{g,i}\in\mathbb{R}^{\hat{C}}$ using the Mamba backbone. These features are then projected into the language space of the LLM using a learnable projection layer, resulting in $\{c_{g,1},c_{g,2},\dotsc,c_{g,n}\},c_{g,i}\in\mathbb{R}^{E}$. This projection aligns the visual features with the text embeddings used by the LLM, facilitating seamless integration of visual and textual information.
$\bullet$ Residual Calculation.  To guide the large language model to generate more accurate medical reports, in this paper, we measure the difference between the current input X-ray sample and context samples, and term the difference as residual tokens. For each context sample, we assign a disease prompt “With disease” or “Normal” and also take the visual-prompt difference into consideration, i.e.,
$R_{v}^{+}$ $=[v_{g}-c_{g,1}^{+},v_{g}-c_{g,2}^{+},\dotsc,v_{g}-c_{g,n}^{+}]$ (9)
$R_{v}^{-}$ $=[v_{g}-c_{g,1}^{-},v_{g}-c_{g,2}^{-},\dotsc,v_{g}-c_{g,n}^{-}]$ (10)
$R_{t}$ $=[v_{g}-t_{1},v_{g}-t_{2},\dotsc,v_{g}-t_{p}]$ (11)
where $c_{g,i}^{+}$ and $c_{g,i}^{-}$ denote the projected and tokenized features of the $i$-th positive and negative image, respectively. [$\cdot$,$\cdot$ , ⋅] denotes the concatenate operation. $R_{v}^{+}$ and $R_{v}^{-}$ represent the residual tokens for the positive and negative examples, while $R_{t}$ are the residual tokens for the disease prompt. $t_{p}$ is the $p$-th tokenized and projected text token.
$\bullet$ Prompt Construction.  All operations are conducted on high-dimensional language space. Each token, including those in the residuals and the original text, is converted into an embedding vector by the previous step, which ensures all elements of the prompt are represented in a manner that the LLM can effectively understand and process. Therefore, the final prompt for the LLM is constructed by concatenating the residuals and tokenized features as follows:
Prompt $=[R_{t},R_{v}^{-},R_{t},R_{v}^{+},R_{t},T,v_{s},T]$ (12)
Here, $T$ and $v_{s}$ are the tokenized text and vision token representations of the current image, respectively. This structured input is fed into the LLM to generate the final medical report, incorporating both visual and textual context in a coherent and informative manner.
3.3.3 LLM Head & Loss Function
The Large Language Model plays a central role in generating detailed and accurate medical reports from X-ray images. As described in the previous sections, the input to the LLM consists of a composite prompt that includes both visual and textual residuals derived from positive and negative context samples, along with the tokenized text corresponding to the input X-ray image. Once the contextual information has been integrated, the LLM generates the medical report. The generation process involves decoding the embedded and contextually enriched prompt into a coherent and comprehensive text. In our experiments, various LLMs are evaluated to achieve higher performance, including lightweight LLM Qwen1.5 (0.5B, 1.8B, 4B, 7B) [2], the medium-sized Llama2 (7B) [38], Llama3 (8B) [13], and large-scale MedicalGPT (13B) [48]. For even larger LLMs, considering the computational cost, this paper will not consider them for the time being.
To optimize our R2GenCSR framework, we adopt the cross-entropy loss function to measure the difference between the generated medical reports and the ground truth annotations. Specifically, we apply instruction-tuning to the LLM to generate medical reports, maintaining its original auto-regressive training objective. During this process, the LLM is fine-tuned specifically on the tokens of the medical report, guided by the instruction prompt that encapsulates the visual and textual residuals. Our loss function is defined as the negative log-likelihood of the sequence of report tokens. This can be formulated as:
$\mathcal{L}=-\sum_{i=1}^{T}\log p_{\theta}(y_{i}|Prompt,Y_{r,<i})$ (13)
where $\theta$ denotes the trainable parameters, and $T$ is the length of the whole medical report. $y_{i}$ is the token being predicted at the current step i, $Prompt$ is the instruction prompt that includes the residuals and tokenized features, and $Y_{r,<i}$ is the sequence of report tokens before the current prediction token $y_{i}$. The instruction-tuning ensures that the LLM generates a report that aligns with the provided instructions and context, thus producing a coherent and informative medical report.
4 Experiment
4.1 Datasets and Evaluation Metrics
We evaluate the performance of our model on three datasets, including IU-XRay [10], MIMIC-CXR [21], and CheXpert Plus [6] dataset. A brief introduction to these datasets is given below.
$\bullet$ IU-XRay [10] is one of the most widely used publicly accessible medical image datasets in the field of medical report generation, which was released on year 2016. It contains 7,470 images and 3,955 radiology reports, each consisting of four parts: indication, comparison, Findings, and Impression. For a fair comparison, we have used the same dataset partition protocols as R2GenGPT and set the training/test/val for the dataset to 7:1:2.
$\bullet$ MIMIC-CXR [21] is a large publicly available dataset of chest radiographs with free-text radiology reports. These records, comprising 377, 110 radiographic images and 227, 835 radiology reports collected from 65, 379 individuals, span the years 2011-2016 and originate from the Beth Israel Deaconess Medical Center Emergency Department in Boston, MA. For a fair comparison, we used the same dataset partition protocols as R2GenGPT, where 270,790 samples were used to train the model, and another 2,130 and 3,858 samples were used as validation and test sets, respectively.
$\bullet$ CheXpert Plus [6] is a large, newly released, organized med dataset with 223K radiology report-X-ray pairs from 64.7K patients, with each report detailed into 11 sections and X-rays in DICOM format with 47 metadata elements. It’s annotated for 14 chest conditions and patient metadata. We utilize the Findings as our ground truth and randomly partition the dataset into a ratio of 7:1:2, which consists of training, testing, and validation sets with 40,463, 5,780, and 11,562 samples respectively. The split protocols for this dataset will be released for other researchers to reproduce our experiments.
For the X-ray medical report generation, we adopt the widely used four metrics for the evaluation, including CIDEr [41], BLEU  [31], ROUGE-L  [25], and METEOR  [3]. Specifically, CIDEr measures the consensus between the generated captions and multiple reference captions. It evaluates the quality of image captioning by computing the cosine similarity between n-grams in the generated caption and those in the reference captions. BLEU evaluates the quality of machine-generated translations or text summaries by comparing them against reference translations or summaries. It measures the precision of n-grams (usually up to 4-grams) in the generated text compared to the reference texts. ROUGE-L assesses the quality of text summaries or translations by comparing them to reference texts. It focuses on the longest common subsequences between the generated and reference texts, emphasizing recall. METEOR evaluates machine-generated translations or summaries by considering both unigram precision and recall, as well as the alignment between the generated and reference texts. It also incorporates stemming and synonymy matching.
4.2 Implementation Details
In our experiments, the input X-ray image is default resized as $224\times 224$, and the beam search is adopted for the report generation. The beam width is set as 5 and 3 for the IU-Xray and MIMIC-CXR datasets, respectively. The training procedure is conducted on a server with NVIDIA A800 80GB GPUs using a mixed precision. We train the proposed R2GenCSR for 20 and 25 epochs on the MIMIC-CXR and IU-Xray dataset. Mini-batch sizes are 36 and 32 for the MIMIC-CXR and IU-Xray datasets, respectively, both trained at a learning rate of 1e-4. The CheXpert Plus dataset adopts the same training protocol as MIMIC-CXR. More details can be found in our source code.
4.3 Comparison on Public Benchmarks
$\bullet$ Results on IU-Xray dataset. As shown in Table 1, we compare our results to state-of-the-art (SOTA) methods on the IU X-Ray datasets. It is important to note that the R2GenGPT method’s reported results were based on a concatenation of Impression and Findings as the testing ground truth, which we believe is not representative of general scenarios so we re-trained their method using only with Findings. Our method demonstrates competitive performance, achieving a BLEU-4 score of 0.206, which surpasses existing methods, highlighting the effectiveness of our context-guided efficient X-ray medical report generation framework. Additionally, our approach attains in the precision of ROUGE-L, METEOR, and CIDEr scores, which are currently at 0.401, 0.412, and 0.579, respectively. These results, albeit not optimal, still affirm the superiority of our approach in generating precise and coherent medical reports compared to current SOTA methods.
$\bullet$ Results on MIMIC-CXR dataset. As shown in Table 1, we report our results on the large-scale MIMIC-CXR dataset. Our method achieved a BLEU-1 score of 0.420, a BLEU-4 score of 0.136, and a ROUGE-L score of 0.291, indicating its ability to generate precise and contextually relevant medical reports. The CIDEr score of 0.267 suggests that the method produces descriptive reports that closely match the reference summaries, highlighting its practical application in clinical settings. Note that our approach employs a dataset-specific strategy, with R2GenCSR-Llama2 being the model variant optimized for the MIMIC-CXR dataset, ensuring that the model is well-suited to the medical report it processes.
Table 1: Comparison of our model’s performance on the IU-Xray and MIMIC-CXR datasets. The symbol $\dagger$ indicates that we follow the R2Gen annotation using Findings and evaluate with our method because their report modifies the ground truth to an Impression concatenated with Findings. The best result is highlighted in bold, and the second-best result is underlined.
Dataset Methods Year BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE-L METEOR CIDEr
IU X-Ray R2Gen [7] EMNLP 2020 0.470 0.304 0.219 0.165 0.371 0.187 -
R2GenCMN [8] ACL-IJCNLP 2021 0.475 0.309 0.222 0.170 0.375 0.191 -
PPKED [27] CVPR 2021 0.483 0.315 0.224 0.168 0.376 0.187 0.351
AlignTrans [52] MICCAI 2021 0.484 0.313 0.225 0.173 0.379 0.204 -
Clinical-BERT [49] AAAI 2022 0.495 0.330 0.231 0.170 0.376 0.209 0.432
METransformer [46] CVPR 2023 0.483 0.322 0.228 0.172 0.380 0.192 0.435
DCL [24] CVPR 2023 - - - 0.163 0.383 0.193 0.586
Token-Mixer [51] IEEE TMI 2024 0.483 0.338 0.250 0.190 0.402 0.208 0.482
PromptMRG [19] AAAI 2024 0.401 - - 0.098 0.160 0.281 -
BootstrappingLLM [26] AAAI 2024 0.499 0.323 0.238 0.184 0.390 0.208 -
[…]
R2GenCSR-Qwen Ours 0.514 0.351 0.262 0.206 0.401 0.215 0.579
MIMIC-CXR R2Gen [7] EMNLP 2020 0.353 0.218 0.145 0.103 0.277 0.142 -
R2GenCMN [8] ACL-IJCNLP 2021 0.353 0.218 0.148 0.106 0.278 0.142 -
PPKED [27] CVPR 2021 0.360 0.224 0.149 0.106 0.284 0.149 0.237
AlignTrans [52] MICCAI 2021 0.378 0.235 0.156 0.112 0.283 0.158 -
Clinical-BERT [49] AAAI 2022 0.383 0.230 0.151 0.106 0.275 0.144 0.151
METransformer [46] CVPR 2023 0.386 0.250 0.169 0.124 0.291 0.152 0.362
DCL [24] CVPR 2023 - - - 0.109 0.284 0.150 0.281
Token-Mixer [51] IEEE TMI 2024 0.409 0.257 0.175 0.124 0.288 0.158 0.163
PromptMRG [19] AAAI 2024 0.398 - - 0.112 0.268 0.157 -
BootstrappingLLM [26] AAAI 2024 0.402 0.262 0.180 0.128 0.291 0.175 -
[…]
R2GenCSR-Llama2 Ours 0.420 0.268 0.186 0.136 0.291 0.167 0.267
$\bullet$ Results on CheXpert Plus dataset. As shown in Table 2, we re-train three R2Gen series medical report generation models on the recently released CheXpert Plus dataset, including R2Gen [7], R2GenCMN [8], and R2Gen-GPT [45]. We can find that the large language model (Llama2) based R2Gen-GPT achieves better performance on all four evaluation metrics. However, our newly proposed R2GenCSR-Llama2 model still beats the R2Gen-GPT and improves 0.001, 0.005, 0.006, and 0.014 on Bleu-4, ROUGE-L, METEOR, and CIDEr metrics, respectively. These experimental results fully validated the effectiveness of our proposed modules for the X-ray medical report generation task.
Table 2: Comparison on the CheXpert Plus dataset.
Method Bleu-4 ROUGE-L METEOR CIDEr
R2Gen [7] 0.091 0.262 0.131 -
R2GenCMN [8] 0.095 0.262 0.142 -
R2Gen-GPT [45] 0.102 0.267 0.157 0.179
R2GenCSR-Qwen1.5 0.100 0.272 0.137 0.159
R2GenCSR-Llama2 0.103 0.272 0.163 0.193
4.4 Component Analysis
In the general scenario, we conduct a detailed component analysis of our proposed framework on the IU-Xray dataset to investigate the impact of each key module, including VMamba, the utilization of context information, fixed pair strategy, and the performance of different Language Models (Qwen1.5 and Llama2), as shown in the Table 3. The results #1 and #4 indicate that with all the component existent can significantly improve the performance across various evaluation metrics (Bleu-1, Bleu-4, ROUGE-L, and METEOR). Specifically, we use Swin Transformer instead while VMamba absent, by comparing #2 with #8 and #6 with #7 that the VMamba module outperforms the Swin Transformer in Bleu-1, Bleu-4, ROUGE-L, it confirm that Vmamba extracting efficient visual features from X-ray images. When comparing #1 with #2 and #4 with #6, the utilization of context is shown to facilitate the model in producing high-quality reports. Furthermore, compared #1 with #3 and #4 with #5, the fixed pair strategy resulting in improved Bleu-4 and METEOR metrics for both Llama2 and Qwen1.5 backends, is contribute to effectively utilizing both positively and negatively samples. The #7 with #8 comparison between Qwen1.5 and Llama2 language models reveals that the choice of language model does indeed influence the final performance, with Qwen1.5 generally yielding superior results. Additionally, our method is shown to be versatile and capable of generalizing even when used with other language models. Based on these comprehensive experiments, we can conclude that our method is effective and robust in generating X-ray medical reports.
Table 3: Component analysis of the key modules in our framework on IU-Xray dataset.
Index VMamba Context Fixed pair Qwen1.5 Llama2 Bleu-1 Bleu-4 ROUGE-L METEOR
#1 0.514 0.206 0.401 0.215
#2 0.498 0.189 0.387 0.203
#3 0.503 0.195 0.389 0.214
#4 0.483 0.180 0.385 0.219
#5 0.496 0.176 0.387 0.209
#6 0.476 0.175 0.377 0.209
#7 0.454 0.173 0.376 0.210
#8 0.478 0.180 0.384 0.195
4.5 Ablation Study
In this section, we conduct a series of ablation studies to evaluate the impact of various components within our proposed context-guided efficient X-ray medical report generation framework.
$\bullet$ Analysis of different VMamba backbones.  As shown in Table 4, we evaluate the Tiny, Small, and Base versions of VMamba on the IU-Xray dataset. As the scale of the VMamba backbone increases, the performance metrics show a consistent improvement. The Base version of VMamba demonstrates the best overall performance, with improvements over the Small version of 0.009, 0.002, 0.001 and 0.003 in Bleu-1, Bleu-4, ROUGE-L, and METEOR, respectively. It confirms that a larger VMamba backbone can capture more intricate visual features from X-ray images, thereby enhancing the quality of the generated medical reports.
Table 4: Results of different VMamba on IU-Xray dataset.
Method Bleu-1 Bleu-4 ROUGE-L METEOR
Tiny 0.484 0.196 0.398 0.205
Small 0.505 0.204 0.391 0.212
Base 0.514 0.206 0.401 0.215
$\bullet$ Analysis of different context sample retrieve methods.  As shown in Table 5, we assess the performance of three distinct methods: Chexbert [34] is utilized to extract 14 medical observations, we define the images corresponding to reports labeled “No Finding” serving as negative samples, while all other images are classified as positive. Despite its moderate performance, Chexbert provides a baseline for comparison. The Random method randomly retrieves context samples and could slightly improve the generation results. The Keyword method retrieves samples based on keyword matching and is the most effective approach, surpassing the other two methods on all evaluated metrics.
Table 5: Results of different context retrieve methods on IU-Xray.
Method Bleu-1 Bleu-4 ROUGE-L METEOR
Chexbert 0.495 0.195 0.391 0.207
Random 0.502 0.198 0.392 0.217
Keyword 0.514 0.206 0.401 0.215
$\bullet$ Analysis of different stages for feature subtraction.  As illustrated in Table 6, we observe that conducting visual subtraction alone in the large language model embedding space yields better results than subtraction outside of it, with a 0.004 improvement in Bleu-4. When both visual and context-instructed text residuals are considered, there is a slight improvement in the metrics after the LLM projection space, with increases of 0.008, 0.006, 0.009 and 0.002 in Bleu-1, Bleu-4, ROUGE-L, and METEOR, respectively. This analysis demonstrates the importance of feature subtraction at different stages and confirms that our approach effectively enhances the feature representation.
Table 6: Feature subtraction at different stages on the IU-Xray dataset. ‘Before-proj. VR’ denotes the visual residual conducted before projection, ‘After-proj. VR’ denotes the visual residual conducted after projection, and ‘After-proj. VR + TR’ denotes the visual and context-instructed text residuals conducted after projection.
Method Bleu-1 Bleu-4 ROUGE-L METEOR
Before-proj. VR 0.503 0.196 0.390 0.211
After-proj. VR 0.506 0.200 0.392 0.213
After-proj. VR + TR 0.514 0.206 0.401 0.215
$\bullet$ Analysis of different resolutions for report generation.  As shown in Table 7, we conducted experiments on the IU-Xray dataset using three image resolutions: 512 $\times$ 512, 448 $\times$ 448, and 224 $\times$ 224. The performance metrics, including Bleu-1, Bleu-4, ROUGE-L, and METEOR, improved by 0.014, 0.010, 0.013 and 0.010 as the resolution decreased from 512 $\times$ 512 to 224 $\times$ 224, respectively. This unexpected trend can be partially explained by the fact that we utilized the Vmamba model which was pre-trained on 224 $\times$ 224 resolution and it is adept at processing and extracting meaningful features from images of this specific resolution.
Table 7: Results under different resolutions on IU-Xray dataset.
Scale Bleu-1 Bleu-4 ROUGE-L METEOR
512 $\times$ 512 0.500 0.196 0.388 0.205
448 $\times$ 448 0.503 0.201 0.385 0.209
224 $\times$ 224 0.514 0.206 0.401 0.215
Qwen1.5-0.5BQwen1.5-1.8BQwen1.5-4BQwen1.5-7BLlama2-7BLlama3-8B MedicalGPT (Llama-13B) 0.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.5scoreEvaluation Metrics vs. LLMBleu-1Bleu-2Bleu-3Bleu-4ROUGE-LMETEOR
Figure 3: Ablation of language model on IU-Xray dataset.
Table 8: Ablation of Positive-Negative and Instruction prompt on IU-Xray dataset.
Method Bleu-1 Bleu-4 ROUGE-L METEOR
[…]
0.514 0.206 0.401 0.215
[…]
0.503 0.196 0.384 0.209
[…]
0.502 0.200 0.393 0.211
[…]
0.499 0.202 0.393 0.207
[…]
0.497 0.199 0.392 0.209
[…]
0.504 0.196 0.393 0.217
[…]
 0.502 0.203 0.388 0.213
$\bullet$ Analysis of different LLMs for report generation.  As illustrated in Figure 3, we present the performance among models of varying sizes and architectures, including the Qwen and Llama series models, Specifically, the Qwen1.5-1.8B model demonstrates a notable improvement in metrics such as Bleu-4 and ROUGE-L by 0.206 and 0.401, respectively. However, the Llama3-8B model, despite its larger size, underperforms compared to the Qwen1.5-4B model, even worse than Llama2-7B. The specialized MedicalGPT (Llama-13B) model exhibits competitive performance within the Llama series, showcasing the potential benefits of domain-specific fine-tuning for large language models in medical report generation tasks. Therefore, selecting an appropriate LLM that balances size, architecture, and domain specialization is necessary for optimal report generation performance.
$\bullet$ Analysis of different number of context samples.  The Table 9 illustrates the performance across different numbers of context sample pairs. The results reveal an optimal point at 3 context sample pairs, which achieves higher scores than 10 pairs, with improvements for Bleu-1, Bleu-4, ROUGE-L, and METEOR of 0.023, 0.024, 0.018, and 0.010, respectively. As the number of context samples increases (from 3 to 10), there is a noticeable decline in performance across all metrics. We suggest that while additional context samples can contribute to a richer feature representation, an excessive number may introduce noise, detracting from the model’s ability to generate accurate and coherent reports. The use of a single context sample pair also yields competitive results, but the slight improvement seen with three pairs indicates that a moderate amount of context can enhance the learning process.
Table 9: Ablation of context samples on IU-Xray dataset.
#Pairs Bleu-1 Bleu-4 ROUGE-L METEOR
1 0.513 0.196 0.390 0.215
3 0.514 0.206 0.401 0.215
5 0.490 0.182 0.375 0.203
10 0.491 0.182 0.383 0.205
$\bullet$ Analysis of VMamba and Swin-Transformer for report generation.  In Table 10, we compare the training and testing efficiency of VMamba and Swin Transformer as vision backbones on a single A800 80G GPU. Although R2GenCSR-VMamba has slightly more trainable parameters (91.7 M vs. 90.9 M) and consumes more memory (75723 MB vs. 70203 MB), its training time requires only 3.98 hours per epoch, which is less than the 5.85 hours per epoch needed for R2GenCSR-Swin. Despite similar testing times for both models, VMamba has a lower FLOPs count (1852.35 G) than Swin Transformer (1855.02 G) and the overall efficiency of R2GenCSR-VMamba is superior.
Table 10: Evaluation between VMamba and Swin Transformer on MIMIC-CXR dataset.
Scale and Efficiency R2GenCSR-VMamba R2GenCSR-SwinT
Trainable Parameter 91.7 M 90.9 M
FLOPs 1852.35 G 1855.02 G
Training Time 3.98 h/Epoch 5.85 h/Epoch
Testing Time 0.52 h 0.52 h
Memory 75723 MB 70203 MB
Figure 4: X-ray image and feature map and its corresponding report on the MIMIC-CXR dataset.
$\bullet$ Analysis of Positive-Negative context-instructed text and Instruction prompt.  As shown in Table 8, we examine the influence of positive-negative context-instructed text and different instruction prompts on our model’s report generation quality. For instance, by modifying the context-instructed text from Note: <Img>$V_{G}$</Img> normal. Note: <Img>$V_{G}$</Img> with disease. to Observation: <Img>$V_{G}$</Img> appears to be normal and healthy. Observation: <Img>$V_{G}$</Img> shows clear signs of disease., the Bleu-4 score decreased by 0.010. Similarly, by changing the instruction prompt from Generate a comprehensive and detailed diagnosis report for this chest xray image. to Construct a full and methodical diagnostic summary for the chest X-ray displayed., the Bleu-4 score decreased by 0.007. Although the differences in performance between the prompts are relatively small, optimizing prompts can subtly affect the model’s output.
4.6 Visualization
$\bullet$ Feature Maps.  The X-ray image and its corresponding feature map are displayed side by side in Figure 4. It is evident from the feature map that the VMamba vision backbone effectively extracts discriminative visual features from the X-ray image, emphasizing regions of interest such as lesions, organs, and other anomalies. These extracted features supply detailed information for subsequent processing.
$\bullet$ X-ray Medical Report.  Figure 5 illustrates a side-by-side comparison of X-ray images alongside their respective ground truths and the reports generated by our model. It is clear that our method is capable of producing reports that closely align with the ground truth, with only minor discrepancies highlighted in mismatching terms but the overall performance of our framework on the MIMIC-CXR dataset is promising.
Image Ground Truth Ours
Refer to caption Pa and lateral views of the chest provided. There is no focal consolidation effusion or pneumothorax. The cardiomediastinal silhouette is normal. Imaged osseous structures are intact. No free air below the right hemidiaphragm is seen. Pa and lateral views of the chest provided. There is no focal consolidation effusion or pneumothorax. The cardiomediastinal silhouette is normal. Imaged osseous structures are intact. No free air below the right hemidiaphragm is seen. Surgical clips project over the right upper quadrant of the abdomen.
Refer to caption Frontal and lateral views of the chest are obtained. the lungs remain hyperinflated suggesting chronic obstructive pulmonary disease. No focal consolidation pleural effusion or evidence of pneumothorax is seen. The cardiac and mediastinal silhouettes are stable and unremarkable. Hilar contours are also stable. Frontal and lateral views of the chest were obtained. The lungs are hyperinflated with flattening of the diaphragms suggesting chronic obstructive pulmonary disease. No focal consolidation pleural effusion or evidence of pneumothorax is seen. The cardiac and mediastinal silhouettes are stable and unremarkable. Bony structures are intact.
Refer to caption As compared to the previous radiograph the patient has received a new dobbhoff tube. The tip of the tube projects over the middle parts of the stomach. the course of the tube is unremarkable. There is no evidence of complications notably no pneumothorax. Otherwise the radiographic appearance of the thoracic organs is similar to the previous examination. As compared to the previous radiograph the patient has received a new nasogastric tube. The course of the tube is unremarkable the tip of the tube projects over the middle parts of the stomach. There is no evidence of complications notably no pneumothorax. Moderate cardiomegaly with bilateral pleural effusions and areas of atelectasis at the lung bases.
Figure 5: X-ray image and its corresponding ground truth, along with the output of our model generation report on the MIMIC-CXR dataset. The mismatch sentence in the reports are highlighted using different colors.
4.7 Limitation Analysis
Although our proposed R2GenCSR achieves better performance on three large-scale report generation benchmark datasets, however, our model may still limited by the following issues: 1). We adopt a simple retrieval strategy for context sample mining, more advanced retrieval techniques can be exploited to achieve better performance; 2). The knowledge about the disease is ignored in our R2GenCSR framework, this information may be useful to guide the X-ray image report generation task. In our future works, we will consider to improve the proposed framework from the aforementioned two aspects.
5 Conclusion
In this paper, we have developed a novel context-guided efficient X-ray medical report generation framework with the aim of enhancing the performance of Large Language Models in clinical settings. We introduced the Mamba model as a linear complexity vision backbone to effectively extract visual features from X-ray images, ensuring that the computational burden is significantly reduced while maintaining performance parity with the robust Transformer architecture. Our approach was complemented by the integration of vision tokens, context information, and prompt statements to facilitate the generation of high-quality medical reports by the LLM. The proposed framework, extensively evaluated on the IU-Xray, MIMIC-CXR, and CheXpert Plus datasets, has demonstrated its efficacy and state-of-the-art performance. Our work not only contributes to the advancement of automated medical report generation but also provides valuable insights for leveraging LLMs in other domains.